Clearview Software, a US exporter, is concerned about the depreciation of JPY against USD due to JPY receivables of JPY800,000,000 in 180 days. To hedge (protect himself/herself) the position, exporter decides to use a combination of the forward and options markets. The Clearview receives the following outright 180-day forward quote from its banker:
USD/JPY 118.40-118.90
Clearview sells JPY 400,000,000 to deliver in 180 days.
To cover JPY 300m portion of the receivable, the company treasurer buys an over the counter put option at a strike price of JPY120, at 2% (2% is the premium) and leaves JPY100m portion of the exposure uncovered.
At the time the option was purchased, the spot rate was USD/JPY117. Clearview's cost of capital is 10%. What is the effective amount of dollars Clearview will get in 180 days if the spot rate on that date is USD/JPY122?
At the time the option was purchased, the spot rate was USD/JPY117. Clearview's cost of capital is 10%. What is the effective amount of dollars Clearview will get in 180 days if the spot rate on that date is USD/JPY122?
Answer #1

Part 1). Forward hedge for 400 million Yen will get 400*10^6/forward rate = 400*10^6/118.40 = 3,378,378.38 USD after 180 days.

Part 2). Option hedging of 300 million Yen:

Option premium (in USD) = (premium*amount)/spot rate = (2%*300*10^6)/117 = 51,282.05 USD

This amount carried forward for 180 days @ 10% will be 51,282.50*(1+ 10%*180/360) = 53,846.15 USD

Since spot rate after 180 days is 122 which is higher than the strike price of 120, the put option will be exercised.

The amount received will be 300*10^6/120 = 2,500,000 USD

Net proceeds = amount recieved - premium = 2,500,000 = 53,846.15 = 2,446,153.85 USD

Part 3). 100 million Yen which was exposed will be sold at the spot rate of 122 after 180 days so amount received will be 100*10^6/122 = 819,672.13 USD

Effective amount (in USD) after 180 days = 3,378,378.38 + 2,446,153.85 + 819,672.13 = 6,644,204.36 USD